what is the slope-intercept of -3/4 and y-intercept of -5
Found 2 solutions by fractalier, Alan3354:
Answer by fractalier(6550)

I'm assuming you are given the slope as -3/4?
If so, we plug into y = mx + b form to get
y = (-3/4)x - 5

Answer by Alan3354(69443)

what is the slope-intercept of -3/4 and y-intercept of -5
----------
Slope m = -3/4
y-intercept b = -5
----
Slope-intercept form is:
y = mx + b